Owl Shift 21-Nov-1999

ERWIN@wishpa.physics.wisc.edu
Sun, 21 Nov 99 07:13 CDT

Shift Summary for hhh shift, dd/mmm/99

On Shift: V. Jerjer, Erwin

Hours of beam available at all intensities: 4.5
Hours of beam available at nominal intensity: 4.0
Hours of E832 data (special runs not included): 4.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Linac problems

========================================================================
Runs ended this shift: 15022, 15023 (Cosmic)
Run 15024 BEAM_799 in progress

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: About 1.5 hour access. Valerie worked on
V0, V1 latches with only a little success. Sasha G. came in to
reconfigure the aerogel test behind the iron.

========================================================================
Overview of shift: Except for Linac failure there was smooth data
taking. Intensity seldom reached the requested 9.5E12, but oscillated
around 9E12. For several hours there were many alarms due to low
live time. These became less frequent towards the end of the shift.
The alarms were correlated with the beam-spill display, but not
in a meaningful way.

========================================================================
Description of Shift:

run spill time comment

------------------------------------------------------------------------
00:00 Shift Change
15022 00:25 Run ended when Linac failed
15023 01:45 Cosmic Run started so Valerie could study V-V' latches
on access. Sasha G. was called so he could come
on access to remove the aerogel from test setup
in order to study wave shifter background.
15024 0 03:08 Beam returns. Start new run.
5:30 Split tape UBH29 mounted.

========================================================================
Follow up Notes on Previous Shift

DC4X Left and Right hole counters were sometimes in disagreement
on the previous shift. No such disagreement occured on this shift.
These hole counters are not presently in the hyperon trigger, which
leads to higher level one rates.

========================================================================
Items to Follow up on Next Shift

Keep an eye on the DC4X Left and Right agreement on the scaler screen.
If there is a disagreement that persists, call in the hyperoners to
search for the problem in upstairs level 1 trigger logic.

The TDC timing for RC7 seems threshold dependent. Plot in the logbook
on p. 270